# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ 3.2 Proof of Theorem 1
It follows by applying [3, Proposition 4.4 (i) and Theorem 4.5 (i)] (with $a^{\prime}=-1$, $a=z_{0}$ and $F(z)\coloneqq\log(1+z)$, for $z\in(-1,\infty)$), Lemma 3 and [8, Lemma 3, Chapter 3, Section 16] that the family $(Z^{(z)})_{z\in\mathbb{N}_{0}}$ of BPI process converges weakly, in $\mathbb{D}(\mathbb{R}_{+},\overline{\mathbb{R}})$, as $z\rightarrow\infty$, towards $Z^{(\infty)}=(Z^{(\infty)}_{t})_{t\geq 0}$. Moreover,
@@
$\mathbb{P}\Big{(}Z^{(\infty)}_{t}<\infty\,\,\text{for all}\,\,t>0\Big{)}=1\qquad\text{and}\qquad\lim_{t\downarrow 0}\frac{Z_{t}^{(\infty)}}{v_{t}}=1,\quad\mathbb{P}-\text{almost surely}.$ (42)
@@
The condition in [3, Proposition 4.4 ] is verified by using Lemma 1 (Recall that, under the assumptions that $\mathbf{m}_{\rm int}<0$ and $\mathbf{m}_{\rm br}<\infty$, the BPI process does not explode in finite time [7, Example 1]). The conditions in [3, Theorem 4.5] are verified by using Lemma 2, Lemma 5 and Lemma 3. Finally, (42) and Lemma 4 imply that $\lim_{t\downarrow 0}tZ_{t}^{(\infty)}=(-\mathbf{m}_{\rm int})^{-1}$, $\mathbb{P}$-almost surely, which concludes our proof. ∎
3.3 Proof of Theorem 2
We now proceed with the proof of Theorem 2. In order to do so, let us deduce some technical Lemmas which will be useful for the sequel.
Lemma 6.
Suppose that $\mathbf{m}_{\rm int}=0$ and (${\bf H2}$) hold. Suppose also that either $\bm{\sigma}^{2}_{{\rm int}}$ is finite or (${\bf H3}$) is satisfied. Then, (i), (ii), (iii) and (iv) in Lemma 2 hold.
We start by proving part of Lemma 6, that is, (i), (ii) and (iv) in Lemma 2 hold under the assumptions of Lemma 6.
Proof of (i), (ii) and (iv) in Lemma 6.
We first observe that the proofs of (i), (ii), and (iv) in Lemma 2 rely solely on the conditions $\sum_{i\geq 1}ib_{i}<\infty$ and $\mathbf{m}_{\rm br}<\infty$. In our setting, the latter follows from (${\bf H2}$) and the former from our assumption that $\mathbf{m}_{\rm int}=0$. ∎
The proof that (iii) in Lemma 2 holds under the assumption of Lemma 6 relies in the following two technical Lemmas. Recall the definition of the function $h$ in Lemma 2 part (ii). This function is well-defined when $\mathbf{m}_{\rm int}=0$ and (${\bf H2}$) holds. Moreover, by (26), $h(z)=h_{\rm br}(z)+h_{\rm int}(z)$, for $z\in[0,\infty)$, where $h_{\rm br}$ and $h_{\rm int}$ are defined in (27) and (28), respectively.
Lemma 7.
Suppose that $\mathbf{m}_{\rm int}=0$, $\bm{\sigma}^{2}_{{\rm int}}<\infty$ and (${\bf H2}$) holds. Then,
$\lim_{z\rightarrow\infty}h(z)=\mathbf{m}_{\rm br}-\frac{\bm{\sigma}^{2}_{{\rm int}}}{2}.$
@@ 3.3 Proof of Theorem 2
Note that (27), (24), (${\bf H2}$) and the dominated convergence imply that
$\lim_{z\rightarrow\infty}h_{\rm br}(z)=\mathbf{m}_{\rm br}.$ (43)
On the one hand, we observe from (28) and (${\bf H2}$), that, for $z\in[0,\infty)$,
$h_{\rm int}(z)$ $=z(z-1)\sum_{i\geq 1}b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}.$ (44)
On the other hand, it follows from (30) and the triangle inequality that, for $z\in(1,\infty)$,
$\Big{|}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{|}$ $=\Big{|}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\frac{i}{z-1}+\log\Big{(}1+\frac{i}{z+1}\Big{)}-\frac{i}{z-1}\Big{|}$
$=\Big{|}i\int_{0}^{-1}\Big{(}\frac{1}{z+1+y}-\frac{1}{z-1}\Big{)}{\rm d}y+\int_{0}^{i}\Big{(}\frac{1}{z+1+y}-\frac{1}{z-1}\Big{)}{\rm d}y\Big{|}$
$\leq\frac{3ib_{i}}{z(z-1)}+\frac{3i^{2}b_{i}}{(z+1)(z-1)}$
and thus,
$\Big{|}z(z-1)b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}\Big{|}\leq 3ib_{i}+3i^{2}b_{i}.$ (45)
Since
$\lim_{z\rightarrow\infty}z(z-1)b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}=-\frac{(i+i^{2})b_{i}}{2},$
we deduce from (44), (45) and the dominated convergence theorem that $\lim_{z\rightarrow\infty}h_{\rm int}(z)=-\frac{\bm{\sigma}^{2}_{{\rm int}}}{2}$. This together with (43) concludes our proof. ∎
Lemma 8.
Suppose that $\mathbf{m}_{\rm int}=0$, (${\bf H2}$) and (${\bf H3}$) hold. Then, there are constants $z^{\prime},C_{1,\alpha},C_{2,\alpha}\in(0,\infty)$ such that $-C_{1,\alpha}z^{2-\alpha}L(z)\leq h(z)\leq-C_{2,\alpha}z^{2-\alpha}L(z)$, for $z\geq z^{\prime}$.
Proof.
We first deduce the lower-bound. By the inequality $\log(1+x)\geq x-\frac{x^{2}}{2}$, for $x\geq 2$, and (24), we have that, for $z\geq 2$,
$\sum_{i=1}^{\lfloor z\rfloor}b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}$ $\geq\sum_{i=1}^{\lfloor z\rfloor}b_{i}\Big{(}-\frac{i}{z}+\frac{i}{z+1}-\frac{i^{2}}{2(z+1)^{2}}\Big{)}$
$=\frac{1}{z(z+1)}\sum_{i=1}^{\lfloor z\rfloor}ib_{i}-\frac{1}{2(z+1)^{2}}\sum_{i=1}^{\lfloor z\rfloor}i^{2}b_{i}.$ (46)
By (24), we also have that
$\sum_{i=\lfloor z\rfloor+1}^{\infty}b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}$ $\geq-\frac{1}{z}\sum_{i=\lfloor z\rfloor+1}^{\infty}ib_{i}.$ (47)
Then, (44), (3.3) and (47) imply that, for $z\geq 2$,
$h_{\rm int}(z)$ $\geq-\frac{z(z-1)}{2(z+1)^{2}}\sum_{i=1}^{\lfloor z\rfloor}i^{2}b_{i}-(z-1)\sum_{i=\lfloor z\rfloor+1}^{\infty}ib_{i}.$ (48)
By applying [9, (4.16) in Lemma 4.4], (${\bf H3}$) implies that
$\sum_{i=1}^{\lfloor z\rfloor}i^{2}b_{i}\sim\frac{z^{2-\alpha}}{2-\alpha}L(z)\quad\text{as}\quad z\rightarrow\infty,$ (49)
and while by applying [9, (4.18) in Lemma 4.4], (${\bf H3}$) also implies that
$\sum_{i=\lfloor z\rfloor+1}^{\infty}ib_{i}\sim\frac{z^{1-\alpha}}{\alpha-1}L(z),\quad\text{as}\quad z\rightarrow\infty.$ (50)
Thus, the lower bound for $h$ follows from (48), (49), (50) and (43).
Now, we prove the upper bound. By the following inequality $\log(1+x)\geq\frac{x}{1+x}$, for $x\geq 0$, and (24), we observe that, for $z\geq 0$,
$\sum_{i=\lfloor z/4\rfloor+1}^{\infty}b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}$ $\leq 0.$ (51)
A Taylor expansion argument allow us to deduce that, for $i=1,\dots,\lfloor z/4\rfloor$,
$\log\Big{(}1+\frac{i}{z+1}\Big{)}$ $=\frac{i}{z+1}-\frac{i^{2}}{2(z+1)^{2}}-\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}\Big{(}\frac{i}{z+1}\Big{)}^{k}$
$\leq\frac{i}{z+1}-\frac{i^{2}}{2(z+1)^{2}}+\frac{1}{3}\Big{(}\frac{i}{z+1}\Big{)}^{3}\sum_{k=3}^{\infty}\Big{(}\frac{i}{z+1}\Big{)}^{k-3}$
$\leq\frac{i}{z+1}-\frac{i^{2}}{2(z+1)^{2}}+\frac{4}{9}\Big{(}\frac{i}{z+1}\Big{)}^{3}\frac{z+1}{z+4/3}.$ (52)
Once again the inequality $\log(1+x)\geq\frac{x}{1+x}$, for $x\geq 0$, and (3.3) imply that, for $z\geq 0$,
$\sum_{i=1}^{\lfloor z/4\rfloor}b_{i}\Big{(}-i\log\Big{(}1+\frac{1}{z}\Big{)}+\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}$ $\leq-\frac{7}{18(z+1)^{2}}\sum_{i=1}^{\lfloor z/4\rfloor}i^{2}b_{i}.$ (53)
Hence from (44), (51) and (53) we deduce, for $z\geq 0$,
$h_{\rm int}(z)$ $\leq-\frac{7z(z-1)}{18(z+1)^{2}}\sum_{i=1}^{\lfloor z/4\rfloor}i^{2}b_{i}$ (54)
and the upper bound follows from (49) and (43). This completes the proof. ∎
With these two results in hand, we can now finish the proof of Lemma 6.
Proof of (iii) in Lemma 6.
If $\mathbf{m}_{\rm int}=0$, then Lemma 7 implies (iii) in Lemma 2 (recall that, by (${\bf H2}$), $\mathbf{m}_{\rm br}-\bm{\sigma}^{2}_{{\rm int}}/2<0$). If assumption (${\bf H3}$) holds, then Lemma 8 and [9, Lemma 4.1 (1)] (with $\varepsilon\in(0,2-\alpha)$) imply that $\lim_{z\rightarrow\infty}h(z)=-\infty$ and thus (iii) in Lemma 2 follows. ∎
We derive analogous results corresponding to Lemmas 3, 4, and 5. Recall the definition of the function $\varphi$ in Lemma 2 (iii), that is, $\varphi(z)=(z+1)h(z)$, for $z\in[0,\infty)$. Recall also that Lemma 6 allows us to choose $\hat{z}>1$ such that (iii) in Lemma 2 is satisfied and set $D\coloneqq(\hat{z},\infty)$.
Lemma 9.
Suppose $\mathbf{m}_{\rm int}=0$ and that (${\bf H2}$) hold. If $\bm{\sigma}^{2}_{{\rm int}}<\infty$ holds, then for any $z\in D$, $\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}=\infty$. On the other hand, if (${\bf H3}$) holds, then for any $z\in D$, $\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}<\infty$.
Proof.
Lemma 7 implies the first statement, while the second one follows from Lemma 8 and [9, Lemma 4.1 (1)] (with $\varepsilon\in(0,2-\alpha)$). ∎
Recall that $v_{t}$ denotes the flow starting from infinity as in (33), for any $t\in[0,T(\infty))$. If $\mathbf{m}_{\rm int}=0$, (${\bf H2}$) and (${\bf H3}$) are satisfied then, by Lemma 9, we have that $v_{t}$ satisfies (34), for any $t\in(0,T(\infty))$.
Lemma 10.
Suppose $\mathbf{m}_{\rm int}=0$, and that (${\bf H2}$) and (${\bf H3}$) hold. Then, $v_{\cdot}$ also satisfies (i) and (ii) in Lemma 4. Moreover, for any $\varepsilon\in(0,2-\alpha)$ there are constants $t^{\prime},C_{1,\alpha,\varepsilon},C_{2,\alpha,\varepsilon}\in(0,\infty)$ such that $C_{1,\alpha,\varepsilon}t^{-\frac{1}{2-\alpha-\varepsilon}}\leq v_{t}\leq C_{2,\alpha,\varepsilon}t^{-\frac{1}{2-\alpha+\varepsilon}}$, for $t\leq t^{\prime}$.
Proof.
The proof of the first claim follows as in the proof of Lemma 4. Indeed, only the assumptions that $\mathbf{m}_{\rm int}=0$ and (${\bf H2}$) are needed. By (i) in Lemma 4 and Lemma 8, there are constants $t_{1},C_{1,\alpha},C_{2,\alpha}\in(0,\infty)$ such that $-C_{1,\alpha}z^{3-\alpha}L(z)\leq\varphi(z)\leq-C_{2,\alpha}z^{3-\alpha}L(z)$, for $z\geq v_{t_{1}}$. Since $v_{\cdot}$ is decreasing (i.e., (ii) in Lemma 4), then it follows that
$\int_{v_{t}}^{\infty}\frac{{\rm d}x}{C_{1,\alpha}x^{3-\alpha}L(x)}\leq\int_{v_{t}}^{\infty}\frac{{\rm d}x}{-\varphi(x)}\leq\int_{v_{t}}^{\infty}\frac{{\rm d}x}{C_{2,\alpha}x^{3-\alpha}L(x)},$ (55)
for all $t\leq t_{1}$. Then, (55) and [9, Lemma 4.1 (1)] imply that for any $\varepsilon\in(0,2-\alpha)$ there are constants $t_{2},C_{3,\alpha},C_{4,\alpha}\in(0,\infty)$ (possibly depending on $\varepsilon$) such that
$\int_{v_{t}}^{\infty}\frac{{\rm d}x}{C_{3,\alpha}x^{3-\alpha+\varepsilon}}\leq\int_{v_{t}}^{\infty}\frac{{\rm d}x}{-\varphi(x)}\leq\int_{v_{t}}^{\infty}\frac{{\rm d}x}{C_{4,\alpha}x^{3-\alpha-\varepsilon}},$ (56)
for all $t\leq t_{2}$. Finally, our second claim follows from (56) and (34). ∎
Proof of Corollary 1.
Lemma 10 implies that $\lim_{t\downarrow 0}tv_{t}=\infty$ and therefore, our claim follows from Theorem 2 (ii). ∎
Lemma 11.
Suppose $\mathbf{m}_{\rm int}=0$ and that (${\bf H2}$) hold. Suppose also that either $\bm{\sigma}^{2}_{{\rm int}}<\infty$ or (${\bf H3}$) are satisfied. Let $V(\cdot)$ be the function defined in (37). Then, for all $t\in(0,\infty)$ and any $\varepsilon>0$,
$\int_{0}^{t}\hat{V}_{\varepsilon}(\hat{z},s){\rm d}s<\infty,$ (57)
where $\hat{V}_{\varepsilon}(\hat{z},t)=\sup_{z\in\mathbb{N}_{0}\cap\mathfrak{D}_{\varepsilon}(\hat{z},t)}\varepsilon^{-2}V(z)$, with $\mathfrak{D}_{\varepsilon}(\hat{z},t)\coloneqq\{z\in(\hat{z},\infty):\log(1+z)\leq\log(1+v_{t})+\varepsilon\}$.
Proof.
The case when $\bm{\sigma}^{2}_{{\rm int}}<\infty$ is satisfied has been proved in Remark 1. Then, suppose that (${\bf H3}$) is satisfied. It follows that, for any $\delta_{1}\in(1,\alpha)$, $\sum_{i\geq 1}i^{\delta_{1}}b_{i}<\infty$, that is, (${\bf H1}$) is satisfied (with $\alpha=\delta_{1}$); see for example, the argument used after the introduction of condition (${\bf H1}$). Hence a similar argument as in the proof of Lemma 5 shows that there exists a constant $C_{\delta_{1}}>0$ such that $V(z)\leq C_{\delta_{1}}z^{2-\delta_{1}}$. The rest of the proof follows similarly as in the proof of Lemma 5. Indeed, Lemma 10 implies that, for any $\delta_{2}\in(0,2-\alpha)$ and $t>0$, there exists a constant $C>0$ such that $v_{s}\leq Cs^{-\frac{1}{2-\alpha+\delta_{2}}}$ for $s\in[0,t]$. Therefore, for all $s\in[0,t]$ and any $\varepsilon>0$, there exists a constant $C_{\delta_{1},\delta_{2}}>0$ such that $\hat{V}_{\varepsilon}(\hat{z},s)\leq C_{\delta_{1},\delta_{2}}s^{-\frac{2-\delta_{1}}{2-\alpha+\delta_{2}}}$. Note that $0<\delta_{2}<2-\alpha<\alpha$ and thus $0<\alpha-\delta_{2}<\alpha$. Therefore, we can choose $\delta_{1}\in(1,\alpha)$ such that $10<\frac{2-\delta_{1}}{2-\alpha+\delta_{2}}<$ 1 which clearly implies (57). ∎
We have now all the ingredients to prove Theorem 2.
Proof of Theorem 2.
Recall that, under $\mathbf{m}_{\rm int}=0$ and (${\bf H2}$) (which implies $\mathbf{m}_{\rm br}<\infty$), the critical BPI process does not explode in finite time; see [14, Theorem 1] or Lemma 23. It follows by applying [3, Proposition 4.4 (i)], Lemma 1 and [8, Lemma 3, Chapter 3, Section 16] that the family $(Z^{(z)})_{z\in\mathbb{N}_{0}}$ of BPI process converges weakly, in $\mathbb{D}(\mathbb{R}_{+},\bar{\mathbb{R}})$, as $z\rightarrow\infty$, towards $Z^{(\infty)}=(Z^{(\infty)}_{t})_{t\geq 0}$.
The claims (i) and (ii) follows from [3, Theorem 4.5] (with $a^{\prime}=-1$, any $a=\hat{z}\in(1,\infty)$ and $F(z)\coloneqq\log(1+z)$, for $z\in(-1,\infty)$) and Lemma 10. The conditions in [3, Theorem 4.5] are verified by using Lemmas 6, 9 and 11. This concludes the proof. ∎
3.4 Proof of Theorem 3
In this section, we prove Theorem 3. The key ingredient is an improvement of the estimate obtained in [3, Lemma 4.6] (see also [3, Theorem 3.2]). To achieve this, we use a supermartingale inequality (see Lemma 22 in Appendix A) similar to that used in the proof of [6, Theorem 2].
Lemma 12.
Suppose that (${\bf H4}$) holds for some $\theta>e^{128\sqrt{2}}$. Then, for $p\in\mathbb{N}$ such that $p\geq 2$ and $z\in\mathbb{N}_{0}$,
$\int_{\mathcal{X}}\Big{|}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}\Big{|}^{p}q({\rm d}u,{\rm d}y,{\rm d}x)\leq\frac{p!}{(\ln\theta)^{p}}\Big{(}d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i}\Big{)}$ (58)
and
$\int_{\mathcal{X}}\Big{|}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}\Big{|}^{2p}q({\rm d}u,{\rm d}y,{\rm d}x)\leq\frac{2^{p}p!}{(\ln\theta)^{p}}\Big{(}d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i}\Big{)}.$ (59)
Proof.
If $z=0$, then $H(0,u,y,x)=0$, for $u=1,2$, $y\in\mathbb{N}\cup\Delta$ and $x\in\mathbb{N}_{-1}$ and thus, (58) is satisfied. By (17) and (19),
@@ 3.4 Proof of Theorem 3
$\quad\quad\quad+\frac{z(z-1)}{2}\int_{\mathbb{N}_{-1}}\Big{|}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{|}^{p}\nu_{\rm int}({\rm d}x).$ (60)
On the one hand, it follows from (14) and (24) that, for $z\in\mathbb{N}$,
$z\int_{\mathbb{N}_{-1}}\Big{|}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{|}^{p}\nu_{\rm br}({\rm d}x)=zd\Big{(}\log\Big{(}1+\frac{1}{z}\Big{)}\Big{)}^{p}+z\sum_{i\geq 1}\pi_{i}\Big{(}\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}^{p}\leq d+\sum_{i\geq 1}i^{p}\pi_{i}.$ (61)
On the other hand, it follows from (14) and (24) that, for $z\in\mathbb{N}$,
@@
$\leq c+\sum_{i\geq 1}i^{p}b_{i}.$ (62)
Note that, for $\theta>e^{128\sqrt{2}}$ as in (${\bf H4}$), we have that $x^{p}\leq\frac{p!}{(\ln\theta)^{p}}\theta^{x}$, for $x\in\mathbb{N}$. Then (58) follows from (3.4), (61), (3.4) and the above inequality.
The proof of (59) is similar, but uses the inequality $\log(1+x)\leq\sqrt{2x}$, for $x\geq 0$, instead of (24). ∎
For the next result, we take $\kappa\geq 0$ as in Lemma 2 (iv), $\theta>e^{128\sqrt{2}}$ as in (${\bf H4}$) and set
$\eta_{\theta}\coloneqq\frac{\ln\theta}{32\sqrt{2}(d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i})}.$
Lemma 13.
@@
$\mathbb{P}\left(\sup_{u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(\infty)}+1}{v_{u}+1}\Big{)}\Big{|}\geq\varepsilon\right)\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$
@@
$T_{D,\varepsilon}(z):=\sup\left\{t\in[0,T(z)):\forall\,s\leq t,\,\overline{B}_{d_{\log}}(\phi(z,s),\varepsilon)\cap\mathbb{N}_{0}\subset D\right\}.$
@@
$\mathbb{P}\left(\sup_{u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(z)}+1}{\phi(z,u)+1}\Big{)}\Big{|}\geq\varepsilon\right)\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$ (63)
@@
$\limsup_{z\rightarrow\infty,z\in\mathbb{N}_{0}}\mathbb{P}\Big{(}\sup_{u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(z)}+1}{\phi(z,u)+1}\Big{)}\Big{|}\geq\varepsilon\Big{)}\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$ (64)
Now, we also fix $t_{0}\in(0,T_{0})$. By Lemma 3, the flow $\phi$ comes down from infinity instantaneously, so $v_{u}<\infty$ on $[t_{0},t]$. By Dini’s theorem, $\phi(z,\cdot)$ converges to $v_{\cdot}$ uniformly on $[t_{0},T_{0}]$, as $z\uparrow\infty$, using the monotonicity of the convergence and the continuity of the limit. We obtain from (64) that for any $t<T_{0}$,
$\limsup_{z\rightarrow\infty,z\in\mathbb{N}_{0}}\mathbb{P}\Big{(}\sup_{t_{0}\leq u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(z)}+1}{v_{u}+1}\Big{)}\Big{|}\geq 2\varepsilon\Big{)}\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}$
and the weak convergence of $(Z^{(z)})_{z\in\mathbb{N}_{0}}$ to $Z^{(\infty)}$ (Theorem 1) yields
$\mathbb{P}\Big{(}\sup_{t_{0}\leq u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(\infty)}+1}{v_{u}+1}\Big{)}\Big{|}\geq 2\varepsilon\Big{)}\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$

### v2

@@ 3.1 Integral representation of BPI processes
It follows by applying Proposition 4.4 (i) and Theorem 4.5 (i) in [3] (with $a^{\prime}=-1$, $a=z_{0}$ and $F(z)\coloneqq\log(1+z)$, for $z\in(-1,\infty)$), Lemma 3.5 and Lemma 3 in Chapter 3, Section 16 of [8] that the family $(Z^{(z)})_{z\in\mathbb{N}_{0}}$ of BPI process converges weakly, in $\mathbb{D}(\mathbb{R}_{+},\overline{\mathbb{R}})$, as $z\rightarrow\infty$, towards $Z^{(\infty)}=(Z^{(\infty)}_{t})_{t\geq 0}$. Moreover,
@@
$\mathbb{P}\Big{(}Z^{(\infty)}_{t}<\infty\,\,\text{for all}\,\,t>0\Big{)}=1\qquad\text{and}\qquad\lim_{t\downarrow 0}\frac{Z_{t}^{(\infty)}}{v_{t}}=1,\quad\mathbb{P}-\text{almost surely}.$ (3.34)
@@
The condition in Proposition 4.4 of [3] is verified by using Lemma 3.3 (Recall that, under the assumptions that $\mathbf{m}_{\rm int}<0$ and $\mathbf{m}_{\rm br}<\infty$, the BPI process does not explode in finite time, see Example 2.1 in [7]). The conditions in Theorem 4.5 of [3] are verified by using Lemma 3.4, Lemma 3.7 and Lemma 3.5. Finally, (3.34) and Lemma 3.6 imply that $\lim_{t\downarrow 0}tZ_{t}^{(\infty)}=(-\mathbf{m}_{\rm int})^{-1}$, $\mathbb{P}$-almost surely, which concludes our proof. ∎
3.3 Proof of Theorem 2.2
In this section, we prove Theorem 2.2. The key ingredient is an improvement of the estimate obtained in Lemma 4.6 of [3] (see also Theorem 3.2 of [3]). To achieve this, we use a supermartingale inequality (see Lemma A.1 in Appendix A) similar to that used in the proof of Theorem 2 of [6].
Lemma 3.9.
Suppose that ($\bf H4$}) hold. For $\theta>1$, $z\in\mathbb{N}_{0}$ and $p\in\mathbb{N}$ such that $p\geq 2$, we have that
$\int_{\mathcal{X}}\Big{|}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}\Big{|}^{p}q({\rm d}u,{\rm d}y,{\rm d}x)\leq\frac{p!}{(\ln\theta)^{p}}\Big{(}d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i}\Big{)}$ (3.35)
and
$\int_{\mathcal{X}}\Big{|}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}\Big{|}^{2p}q({\rm d}u,{\rm d}y,{\rm d}x)\leq\frac{2^{p}p!}{(\ln\theta)^{p}}\Big{(}d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i}\Big{)}.$ (3.36)
@@
If $z=0$, then $H(0,u,y,x)=0$, for $u=1,2$, $y\in\mathbb{N}\cup\Delta$ and $x\in\mathbb{N}_{-1}$ and thus, (3.35) is satisfied. By (3.4) and (3.6),
@@
$\quad\quad\quad+\frac{z(z-1)}{2}\int_{\mathbb{N}_{-1}}\Big{|}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{|}^{p}\nu_{\rm int}({\rm d}x).$ (3.37)
On the one hand, it follows from (3.1) and (3.14) that, for $z\in\mathbb{N}$,
$z\int_{\mathbb{N}_{-1}}\Big{|}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{|}^{p}\nu_{\rm br}({\rm d}x)$ $=zd\Big{(}\log\Big{(}1+\frac{1}{z}\Big{)}\Big{)}^{p}+z\sum_{i\geq 1}\pi_{i}\Big{(}\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}^{p}$
$\leq d+\sum_{i\geq 1}i^{p}\pi_{i}.$ (3.38)
On the other hand, it follows from (3.1) and (3.14) that, for $z\in\mathbb{N}$,
@@
$\leq c+\sum_{i\geq 1}i^{p}b_{i}.$ (3.39)
Note that, for $\theta>1$, we have that $x^{p}\leq\frac{p!}{(\ln\theta)^{p}}\theta^{x}$, for $x\in\mathbb{N}$. Then (3.35) follows from ($\bf H4$}), (3.3), (3.3), (3.3) and the above inequality.
The proof of (3.36) is similar, but uses the inequality $\log(1+x)\leq\sqrt{2x}$, for $x\geq 0$, instead of (3.14). ∎
For the next result, we assume ($\bf H4$}), take $\kappa\geq 0$ as in Lemma 3.4 (iv), $\theta>1$ and set
$\eta_{\theta}\coloneqq\frac{\ln\theta}{32\sqrt{2}(d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i})}.$ (3.40)
Lemma 3.10.
@@
$\mathbb{P}\left(\sup_{u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(\infty)}+1}{v_{u}+1}\Big{)}\Big{|}\geq\varepsilon\right)\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$ (3.41)
@@
$T_{D,\varepsilon}(z):=\sup\left\{t\in[0,T(z)):\forall\,s\leq t,\,\overline{B}_{d_{\log}}(\phi(z,s),\varepsilon)\cap\mathbb{N}_{0}\subset D\right\}.$ (3.42)
@@
$\mathbb{P}\left(\sup_{u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(z)}+1}{\phi(z,u)+1}\Big{)}\Big{|}\geq\varepsilon\right)\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$ (3.43)
@@
$\limsup_{z\rightarrow\infty,z\in\mathbb{N}_{0}}\mathbb{P}\Big{(}\sup_{u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(z)}+1}{\phi(z,u)+1}\Big{)}\Big{|}\geq\varepsilon\Big{)}\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$ (3.44)
Now, we also fix $t_{0}\in(0,T_{0})$. By Lemma 3.5, the flow $\phi$ comes down from infinity instantaneously, so $v_{u}<\infty$ on $t_{0},t]$. By Dini’s theorem, $\phi(z,\cdot)$ converges to $v_{\cdot}$ uniformly on $t_{0},T_{0}]$, as $z\uparrow\infty$, using the monotonicity of the convergence and the continuity of the limit. We obtain from (3.44) that for any $t<T_{0}$,
$\limsup_{z\rightarrow\infty,z\in\mathbb{N}_{0}}\mathbb{P}\Big{(}\sup_{t_{0}\leq u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(z)}+1}{v_{u}+1}\Big{)}\Big{|}\geq 2\varepsilon\Big{)}\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}$ (3.45)
and the weak convergence of $(Z^{(z)})_{z\in\mathbb{N}_{0}}$ to $Z^{(\infty)}$ (Theorem 2.1) yields
$\mathbb{P}\Big{(}\sup_{t_{0}\leq u\leq t}\Big{|}\log\Big{(}\frac{Z_{u}^{(\infty)}+1}{v_{u}+1}\Big{)}\Big{|}\geq 2\varepsilon\Big{)}\leq e^{-\frac{\varepsilon\ln\theta}{64\sqrt{2}}}.$ (3.46)